1. Solve these recurrence relations:

a. f_{n+1} =f_{n}^{2} , Initial condition: f_{0}=2

b. f_{n+1} -f_{n} - n = 0

c. 1+ \sum_{i=0}^{n-1}f_{i} = f_{n+1} , Initial conditions: f_{0} = 1, f_{1} = 1, n\geq 1
Answer #1

1 gruar a) futi itu o-2 but nao 4:40:21 す。( 4 10:1 2 パンツ 2 M す。(12) みた () 400030 b)ne-fu-n=0 but, and う mal 4ーー10 =)二和 , チューチ- fn tn = Fot (or (+3+6+10+15+ = tot EM-) * we dont have initial value given, It Ž fi= turi 1-0 put n=1 Ito=7 = 1/2 =) n=2,